The length of one leg of a right triangle is 18 ft. The length of the hypotenuse is 6ft longer than the other leg. Find the length of the hypotenuse and the other leg.
Answer by midwood_trail(310)

The length of one leg of a right triangle is 18 ft. The length of the hypotenuse is 6ft longer than the other leg. Find the length of the hypotenuse and the other leg.
one leg = 18 ft
hypotenuse = x + 6 ft
the other leg = x
Here is the equation:
x^2 + (18)^2 = (x + 6)^2
Can you take it from here?